Brian and Colin are marking exam papers. Each set takes Brian 40 minutes and Colin 1 hour. Express the times Brian and Colin take as a ratio.
Give your answer in its simplest form.
1
Expert's answer
2020-02-09T16:54:11-0500

In order to solve the problem we need to convert 1 hour into minutes. 1 hour = 60 minutes.

Now we can find a ratio (Brian)/(Colin) = 40(min)/60(min) = [Dividing by their greatest common factor 20] = 2/3.

The answer is 2/3.